# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

### v1

@@ IV-C1 Feature-based Methods
  • Bukhari et al. [15] that takes lexical, syntactic features, and their combination in C source code as the input. This method uses Random Forest, SVM, KNN, and XGBoost to build the detection methods.
IV-C2 Pre-Training-based Method
The only one pre-training-based method is as follows.
  • GPTSniffer [20] that transforms the source code into a sequence of independent tokens. These tokens were processed by the pre-trained model, i.e., CodeBERT. This method uses CodeBERT to develop the detection method.
IV-C3 Perplexity
@@
  • Log-p(x) [21] that utilizes the average token-wise log probability to calculate the perplexity. The log probability is the logarithm of the probability of the next token (ground truth) based on previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a larger log-p(x) than HaCode.
@@
  • Entropy [22] that uses the average entropy to calculate the perplexity. Entropy is the sum of the probability multiplied by the negative logarithm of the probability of each token. This method assumes that LLMgCode has a larger entropy than HaCode.
@@
  • Log-Rank/Rank [22, 23] that utilizes the average log-rank/rank of each token to calculate the perplexity. Log-Rank/Rank is the logarithm of rank/rank of the next token in the vocabulary list. The vocabulary list is sorted according to the probability of each token given the previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a lower log-rank/rank than HaCode.
@@
  • LRR [25] that divides log-likelihood by log-rank to calculate the perplexity. Log-likelihood is the average token-wise log probability, representing the absolute confidence for the true token. Log-rank is the average token-wise log rank, representing the relative confidence for the true token. For LLMgCode, log-rank is more discernible than the log-likelihood. This method assumes that LLMgCode has a larger LRR than HaCode.
@@
  • DetectGPT [24] that leverages the discrepancy between the log probability of the original sample (a collection of tokens) and the average log probability of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by masking and fulfilling some tokens. This method assumes that LLMgCode has a greater perturbation discrepancy as it is more sensitive to perturbations than HaCode.
@@
  • NPR [25] that divides the average log-rank of the perturbed samples by the log-rank of the original sample to calculate the perplexity. The perturbed samples are obtained by replacing some tokens. This method assumes that LLMgCode has a larger NPR as it is more susceptible to perturbations than HaCode.
@@
  • DetectCodeGPT [14] that utilizes the discrepancy between the NPR of the original sample and the average NPR of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by adding spaces and new lines to the original sample. This method assumes that the average NPR of perturbed LLMgCode drops more than HaCode.
IV-D Experimental Settings
@@
Following related work [14, 19, 24, 25], we use the AUC (Area Under the Curve) to measure each detection method. AUC stands for the area under ROC (Receiver Operating Characteristic curve). Given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the calculation of AUC is as follows.
@@
For $\text{RQ}_{1.1}$, we evaluated Perplexity and pre-training-based methods in all the eight programming languages. We evaluated feature-based methods on specific languages because those features are carefully designed for those languages. Due to the time and cost limit, we tested 500 pairs of human-LLM code snippets for each aspect regarding each research question. For $\text{RQ}_{1}$, we tested 500 pairs of human-LLM code snippets for each programming language, difficulty level, and scale. For $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we tested 500 pairs of human-LLM code snippets for each LLM. We randomly selected 500 pairs of human-LLM code snippets from the dataset each time. Following Shi et al. [14], we utilized CodeT5-770M to generate perturbed code snippets for DetectGPT. We added spaces and empty lines to generate perturbed code snippets for NPR and DetectCodeGPT. We generated 50 perturbed samples for each code sample. We carried out the experiments using 1 NVIDIA A100-PCIE-40GB.
V Results and Analysis
@@
V-A $RQ_{\text{1.1}}$: Programming Language
Table I and Table II shows the AUC of Perplexity, and feature-based and pre-training-based methods from the perspective of programming languages, respectively. As the feature-based detection methods are designed for specific languages, we provide only the results in Python (Idialu) and C (Bukhari) in Table II. For presentation clarity, we highlight notable results with blue backgrounds (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
TABLE I: The AUC of Perplexity in different programming languages
Lan Code LLM without perturbation with perturbation
@@
  • 1
    For succinctness, we abbreviate the baseline names as follows: log-p (Log-p(x)), Ent (Entropy), L-Rank (Log-Rank), DGPT (DetectGPT), DCGPT (DetectCodeGPT).
TABLE II: The AUC of feature-based and pre-training-based methods in different programming languages
Lan Feature Pre-training
@@
When comparing Perplexity against the other two classes of methods, we select the best-performing method as representative [^2] For succinctness, we adopt this measure throughout the paper. Perplexity does not perform as well as GPTSniffer which is a pre-training-based method in all the eight programming languages. The AUC of Perplexity and GPTSniffer for each language are as follows: C (90.40%–99.13%), C# (90.12%–98.15%), C++ (91.37%–99.58%), Go (72.68%–94.19%), Java (85.54%–96.11%), Python (63.89%–95.17%), Ruby (73.44%–93.25%), and Rust (89.72%–98.64%). The difference can be explained that GPTSniffer is a supervised method while Perplexity is not. Moreover, Perplexity cannot compare to the feature-based methods in Python and C: Python (63.89%–89.60%), C (90.40%–99.01%).
When comparing the AUC of Perplexity between programming languages, we observe that Perplexity performs best in C/C++ in which NPR and DetectCodeGPT achieve an average AUC of 85%. However, NPR, DetectCodeGPT, and other methods cannot achieve an AUC of over 80% in other languages. Taking C/C++ and Python as examples. C/C++ requires developers to implement the majority of functions manually, which however are provided by third-party libraries in Python. In this case, C/C++ has more code stylometry than Python [64]. However, the design philosophy of Python encourages simplified code stylometry, making it easy for LLMs to generate human-style code.
$\textbf{RQ}_{\textbf{1.1}}$: Perplexity can not perform as well as the feature-based and pre-training-based methods from the perspective of programming languages. Among eight languages, Perplexity is only effective in C and C++.
V-B $RQ_{\text{1.2}}$: Degree of Difficulty
Table III and Table IV present the AUC of Perplexity, and feature-based and pre-training-based methods from the perspective of the degree of difficulty, respectively.
TABLE III: The AUC of Perplexity in different degrees of difficulty
Diff Code LLM without perturbation with perturbation
@@
TABLE IV: The AUC of feature-based and pre-training-based methods in different degrees of difficulty
Diff Feature Pre-training
@@
Perplexity exhibits lower performance compared with the feature-based and pre-training-based methods in terms of all three levels of difficulty. The differences between Perplexity and pre-training-based methods for each level of difficulty are as follows: easy (75.41%–90.57%), middle (77.00%–94.32%), and hard (76.48%–94.33%). When compared with feature-based methods, Perplexity is less effective than Idialu: easy (75.41%–85.37%), middle (77.00%–89.60%), and hard (76.48%–93.18%), and Bukhari: easy (75.41%–96.40%), middle (77.00%–98.98%), and hard (76.48%–98.25%). The discrepancy can be explained that feature-based and pre-training-based methods are based on supervised training while Perplexity not.
When it comes to the AUC of Perplexity between different degrees of difficulty, we find that Perplexity performs not well in all three levels of difficulty, but slightly improves from the easy to hard level. To be specific, log p(x) and Log Rank are the best-performing methods in Perplexity without perturbation, achieving the AUC of 61.25% and 60.10% at the easy level, 62.00%, and 61.76% at the middle level, and 63.54% and 64.53% at the hard level, respectively. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve the AUC of 66.40% and 65.60% at the easy level, 67.82% and 67.67% at the middle level, and 68.95% and 69.64% at the hard level, respectively. Since those solutions to difficult puzzles have complex logic and customized implementations, they have more code stylometry than solutions to easy puzzles.
$\textbf{RQ}_{\textbf{1.2}}$: Perplexity perform not well in all three levels of difficulty, although the performance gradually and slightly improves from the easy to the hard level. The feature-based and pre-training-based methods perform well at all three levels of difficulty.
V-C $RQ_{\text{1.3}}$: Scale of Solution
Table V and Table VI show the AUC of three classes of detection methods, respectively. In Table VI, as there is no large-scale code solution in Python, we provide only the results in the small and middle scale for Idialu.
TABLE V: The AUC of Perplexity in different scales of solution
@@
TABLE VI: The AUC of feature-based and pre-training-based methods in different scales of solution
@@
Perplexity does not perform as well as the pre-training-based method, i.e., GPTSniffer in all the three scales. The AUC of Perplexity and GPTSniffer for each scale are: small (69.75%–82.74%), middle (77.42%–96.10%), and large (87.81%–94.19%). For the featured-based method, Perplexity is less effective than Idialu: small (69.75%–85.87%), middle (77.42%–92.58%), and Bukhari: small (69.75%–91.20%), middle (77.42%–98.26%), and large (87.81%–96.67%). These differences are mainly because feature-based and pre-training-based methods are supervised methods while Perplexity is a collection of heuristic methods.
@@
When comparing the methods in Perplexity regarding different scales, we observe that those well-performed methods improve from the small to large scale. For example, log p(x) is the best-performing method in Perplexity without perturbation, with an AUC of 61.41% for small, 66.97% for middle, and 77.53% for large. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve an AUC of 64.87% and 63.06% for small, 72.91% and 72.19% for middle, and 76.81% and 78.10% for large, respectively. These results can be explained that large-scale code snippets have more diverse and detailed implementations than others, which makes HaCode more distinct from LLMgCode.
@@
$\textbf{RQ}_{\textbf{1.3}}$: Perplexity performs less effectively than the feature-based and pre-training-based methods in all the three scales. Perplexity only performs adequately in the code with large scale.
V-D $RQ_{\text{2}}$: Generalization Capability
Table VII and Table VIII present the AUC of Perplexity and the other two classes of methods in detecting code generated by four LLMs.
TABLE VII: The AUC of Perplexity in different LLMs
Model Code LLM without perturbation with perturbation
log-p Ent Rank L-Rank LRR DGPT NPR DCGPT
GPT-3.5 Incoder 51.89 46.86 39.93 63.01 72.28 49.68 73.10 72.78
Phi-1 75.44 41.93 66.19 75.87 45.27 52.79 81.10 81.64
StarCoder 56.14 54.82 55.86 61.09 57.55 47.53 70.86 70.03
CodeGen2 63.27 49.26 58.56 66.84 56.52 48.27 73.31 72.40
WizCoder 66.34 42.77 60.98 68.54 50.25 72.24 74.26 74.37
CodeLlama 59.41 53.78 60.53 65.79 60.39 41.48 71.52 71.28
Average 62.08 48.24 57.01 66.86 57.04 52.00 74.03 73.75
GPT-4o Incoder 55.29 44.88 44.15 63.16 63.12 54.10 74.89 74.71
Phi-1 72.89 38.02 63.79 72.56 44.10 58.14 80.52 80.88
StarCoder 56.04 52.69 53.07 58.90 53.02 51.63 71.25 70.44
CodeGen2 62.52 48.15 54.12 63.91 51.03 53.03 72.10 71.55
WizCoder 65.54 40.62 57.90 65.91 45.73 66.75 72.95 73.00
CodeLlama 60.70 50.93 56.88 64.68 54.86 48.01 71.13 70.93
Average 62.16 45.88 54.98 64.85 51.98 55.28 73.81 73.59
Gemini Incoder 54.53 45.43 42.55 63.70 76.00 51.37 73.74 73.17
Phi-1 74.55 42.25 71.10 74.30 51.08 54.88 81.78 81.66
StarCoder 60.58 53.99 64.18 64.75 66.84 49.24 73.36 72.26
CodeGen2 67.39 48.34 65.84 69.68 65.25 50.68 76.23 74.98
WizCoder 69.21 44.70 64.43 67.99 50.52 58.07 74.36 74.15
CodeLlama 63.76 53.00 68.01 66.36 68.45 43.83 73.49 73.21
Average 65.00 47.95 62.68 67.80 63.02 51.35 75.49 74.90
Llama Incoder 57.00 42.88 43.15 65.10 74.82 50.84 73.20 73.14
Phi-1 74.45 41.33 68.64 73.47 46.39 54.96 79.42 79.55
StarCoder 66.20 47.63 67.67 70.29 72.05 51.47 75.56 75.23
CodeGen2 71.42 42.79 66.51 73.05 66.92 49.79 77.42 76.95
WizCoder 74.69 38.59 68.99 76.39 56.50 67.27 79.08 78.74
CodeLlama 69.59 47.23 71.41 73.78 72.37 52.15 77.52 77.69
Average 68.89 43.41 64.39 72.01 64.84 54.41 77.03 76.88
TABLE VIII: The AUC of feature-based and pre-training-based methods in different LLMs
LLM Feature Pre-training
Idialu (Python) Bukhari (C) GPTSniffer
GPT-3.5 91.43 98.63 92.21
GPT-4o 83.16 98.49 90.30
Gemini 62.77 92.13 75.21
Llama 71.76 97.15 83.26
Perplexity has a better generalization capacity than the feature-based and pre-training-based methods. The AUC of Perplexity for the four LLMs is as follows: GPT3.5 (81.64%), GPT-4o (80.88%), Gemini (81.78%), and Llama (79.55%), with a range of 2%. The AUC of pre-training-based methods for each LLM is as follows: GPT3.5 (92.21%), GPT-4o (90.30%), Gemini (75.21%), and Llama (83.26%). The max-min difference is over 17%. These differences can be explained that pre-training-based methods are trained on the code generated by GPT-3.5, while Perplexity is LLM-independent. When it comes to feature-based methods, Perplexity outperforms Idialu and Bukhari. The AUC of Idialu for each LLM is as follows: GPT3.5 (91.43%), GPT-4o (83.16%), Gemini (62.77%), and Llama (71.76%). The max-min difference is 29%. the AUC of Bukhari for each LLM is as follows: GPT3.5 (98.63%), GPT-4o (98.49%), Gemini (92.13%), and Llama (97.15%), with a range of 7%. The max-min range of Bukhari is relatively small. Bukhari only works for C in which HaCode is more different from LLMgCode than other languages, making the detection method easier to detect LLMgCode from HaCode.
$\textbf{RQ}_{\textbf{2}}$: Perplexity has a better generalization capability than the feature-based and pre-training-based methods. The AUC of supervised methods drops significantly when training and testing sets are generated by different LLMs.
V-E $RQ_{\text{3}}$: Inference Efficiency

### v2

@@ 4.3.1. Feature-based Methods
  • Bukhari et al. (Bukhari et al., 2023) takes a combination of lexical and syntactic features of C programs as the input. The lexical features are derived from the feature set introduced by Caliskan-Islam (Caliskan-Islam et al., 2015) for code stylometry. The syntactic features are extracted from abstract syntax trees. Bukhari uses Random Forest, SVM, KNN, and XGBoost to build the detection methods. XGBoost outperforms the other classifiers in terms of detection performance.
4.3.2. Pre-Training-based Method
Pre-training-based methods take code snippets as input, convert them into tokens, embed them as vectors, and leverage a pre-trained model for classification. We investigate one pre-training-based method as follows.
  • GPTSniffer (Nguyen et al., 2024) designs rewriting rules (e.g., rewriting class names) to create diverse derivations of the original code samples. These derivations are then used to fine-tune a pre-trained model, i.e., CodeBERT, as the detector. To perform detection, GPTSniffer transforms the input code into a sequence of independent tokens, which are then fed into the fine-tuned CodeBERT to predict whether the code is LLMgCode.
4.3.3. Perplexity
@@
  • Log-p(x) (Solaiman et al., 2019) utilizes the average token-wise log probability to calculate the perplexity. The log probability is the logarithm of the probability of the next token (ground truth) based on previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a larger log-p(x) than HaCode. The calculation of log-p(x) is shown in Equation 1, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
[…]
@@
  • Entropy (Gehrmann et al., 2019) uses the average entropy to calculate the perplexity. Entropy is the sum of the probability multiplied by the negative logarithm of the probability of each token. This method assumes that LLMgCode has a larger entropy than HaCode. The calculation of entropy is shown in Equation 2, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
    (2) $\textrm{Entropy}=-\frac{1}{t}\sum_{i=1}^{t}p_{\theta}(x_{i}\mid x_{<i})\log p_{\theta}(x_{i}\mid x_{<i})$
@@
  • Log-Rank/Rank (Gehrmann et al., 2019; Ippolito et al., 2019) utilizes the average log-rank/rank of each token to calculate the perplexity. Log-Rank/Rank is the logarithm of rank/rank of the next token in the vocabulary list. The vocabulary list is sorted according to the probability of each token given the previous tokens, and the probability is calculated by a language model. This method assumes that LLMgCode has a lower log-rank/rank than HaCode. The calculations of Log-Rank and Rank are shown in Equation 3 and Equation 4 respectively, where $r_{\theta}(x_{i}\mid x_{<i})$ is the rank of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
[…]
    (4) $\textrm{Rank}=\frac{1}{t}\sum_{i=1}^{t}r_{\theta}(x_{i}\mid x_{<i})$
@@
  • LRR (Su et al., 2023) divides log-likelihood by log-rank to calculate the perplexity. Log-likelihood is the average token-wise log probability, representing the absolute confidence for the true token. Log-rank is the average token-wise log rank, representing the relative confidence for the true token. For LLMgCode, log-rank is more discernible than the log-likelihood. This method assumes that LLMgCode has a larger LRR than HaCode. The calculation of LRR is shown in Equation 5, where $p_{\theta}(x_{i}\mid x_{<i})$ is the probability of token $x_{i}$ conditioned on the previous tokens, $r_{\theta}(x_{i}\mid x_{<i})$ is the rank of token $x_{i}$ conditioned on the previous tokens, $t$ is the number of tokens in the sequence.
    (5) $\textrm{LRR}=-\frac{\sum_{i=1}^{t}\log p_{\theta}(x_{i}\mid x_{<i})}{\sum_{i=1}^{t}\log r_{\theta}(x_{i}\mid x_{<i})}$
@@
  • DetectGPT (Mitchell et al., 2023) leverages the discrepancy between the log probability of the original sample (a collection of tokens) and the average log probability of the perturbed samples to calculate the perplexity, normalized by the standard deviation of the log probability of the perturbed samples. The perturbed samples are obtained by masking and completing some tokens. This method assumes that LLMgCode has a greater perturbation discrepancy as it is more sensitive to perturbations than HaCode. The calculation of DetectGPT is shown in Equation 6, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log p_{\theta}(x)$ is the log probability of the original sample $x$, $\log p_{\theta}(\tilde{x}_{i})$ is the log probability of the perturbed sample $\tilde{x}_{i}$, $k$ is the number of perturbed samples.
    (6) $\textrm{DetectGPT}=\frac{\log p_{\theta}(x)-\frac{1}{k}\sum_{i=1}^{k}\log p_{\theta}(\tilde{x_{i}})}{\sqrt{\frac{1}{k-1}\sum_{p=1}^{k}{(\log p_{\theta}(\tilde{x}_{p})-\frac{1}{k}\sum_{i=1}^{k}\log p_{\theta}(\tilde{x}_{i}))}^{2}}}$
@@
  • NPR (Su et al., 2023) divides the average log-rank of the perturbed samples by the log-rank of the original sample to calculate the perplexity. The perturbed samples are obtained by replacing some tokens. This method assumes that LLMgCode has a larger NPR as it is more susceptible to perturbations than HaCode. The calculation of NPR is shown in Equation 7, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log r_{\theta}(x)$ is the log rank of the original sample $x$, $\log r_{\theta}(\tilde{x}_{p})$ is the log rank of the perturbed sample $\tilde{x}_{p}$, $k$ is the number of perturbed samples.
    (7) $\textrm{NPR}=\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(x)}$
@@
  • DetectCodeGPT (Shi et al., 2025) utilizes the discrepancy between the NPR of the original sample and the average NPR of the perturbed samples to calculate the perplexity. The perturbed samples are obtained by adding spaces and new lines to the original sample. This method assumes that the average NPR of perturbed LLMgCode drops more than HaCode. The calculation of NPR is shown in Equation 8, where $x$ is the original sample, $\tilde{x}$ is the perturbed sample, $\log r_{\theta}(x)$ is the log rank of the original sample $x$, $\log r_{\theta}(\tilde{x}_{i})$ is the log rank of the perturbed sample $\tilde{x}_{i}$, $k$ is the number of perturbed samples.
    (8) $\textrm{DetectCodeGPT}=\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(x)}-\frac{1}{k}\sum_{i=1}^{k}\frac{\frac{1}{k}\sum_{p=1}^{k}\log r_{\theta}(\tilde{x}_{p})}{\log r_{\theta}(\tilde{x}_{i})}$
When selecting experimental subjects, we included several detection methods that were not originally proposed to detect LLMgCode, e.g., DetectGPT and NPR. They were proposed to detect LLM-generated text. Since source code is one type of text, these detection methods are also suitable for detecting LLMgCode. This setting was also used in previous related work (Shi et al., 2025; Suh et al., 2025). Although Perplexity is an unsupervised method, we compared it against supervised methods because Shi et al. (2025) reported that Perplexity outperformed supervised methods. We excluded methods whose implementation details are unclear. For example, we excluded AIgCode Detector (Xu and Sheng, 2024) because it did not specify how perturbations are performed.
4.4. Experimental Settings
@@
Following related work (Shi et al., 2025; Idialu et al., 2024; Mitchell et al., 2023; Su et al., 2023), we use the AUC (Area Under the Curve) to evaluate each detection method. AUC stands for the area under ROC (Receiver Operating Characteristic curve). Given a set of TPR (True Positive Rates) and FPR (False Positive Rates) across different thresholds, the calculation of AUC is as follows.
@@
For $\text{RQ}_{1.1}$, we evaluated Perplexity and pre-training-based methods in all eight programming languages. We evaluated feature-based methods on specific languages because the features are designed for those languages only. Due to the time and cost limit, we tested 500 pairs of human-LLM code snippets for each aspect regarding each research question. For $\text{RQ}_{1}$, we tested 500 pairs of human-LLM code snippets for each programming language, difficulty level, and scale. For $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we tested 500 pairs of human-LLM code snippets for each LLM. We randomly selected 500 pairs of human-LLM code snippets from the dataset each time. We selected 500 pairs for the reason of budget and time costs, and such a scale is comparable to the prior related work Shi et al. (2025) (500) and Suh et al. (2025) (400). Following Shi et al. (2025), we utilized CodeT5-770M to generate perturbed code snippets for DetectGPT. We added spaces and empty lines to generate perturbed code snippets for NPR and DetectCodeGPT. We generated 50 perturbed samples for each code sample. Code-specific LLMs are pre-trained using code corpora, which equips them with a deeper understanding of programming languages. We use code-specific LLMs to compute the perplexity of code snippets, as they are more capable of understanding code than general LLMs. Following Shi et al. (2025), we utilized Incoder (1.3B), Phi-1 (1.3B), StarCoder (3B), WizardCoder (3B), CodeGen2 (3.7B), and CodeLlama (7B) to calculate the perplexity for each Perplexity. Table 1 summarizes the specific usage of all LLMs involved in our study. We carried out the experiments using 1 NVIDIA A100-PCIE-40GB.
Table 1. Summary of LLM usage
LLM Usage
GPT3.5 To generate 11,664 LLMgCode solutions for all the evaluation aspects.
GPT-4o To generate 500 LLMgCode solutions for the generalization capability.
Gemini-1.0
Llama-3.1
Incoder (1.3B) To calculate perplexity values for all methods in Perplexity.
Phi-1 (1.3B)
StarCoder (3B)
CodeGen2 (3B)
WizCoder (3.7B)
CodeLlama (7B)
CodeT5 (770M) To generate perturbed code snippets for DetectGPT.
5. Results and Analysis
@@
5.1. $\text{RQ}_{\text{1.1}}$: Programming Language
LABEL:tab:performance-perplexity and Table 3 show the AUC of the three groups of detection methods from the perspective of programming languages. The code LLMs listed in LABEL:tab:performance-perplexity are used to calculate perplexity values. Since feature-based detection methods are designed for specific languages, we provide results only in Python (Idialu) and C (Bukhari) in Table 3. The notable results are highlighted in blue (AUC:$90\%-100\%$;$80\%-90\%$;$70\%-80\%$).
Table 2. The AUC of Perplexity in terms of programming languages
Language Code LLM without perturbation with perturbation
@@
Table 3. The AUC of feature- and pre-training-based methods in terms of programming languages
Language Feature Pre-training
@@
When comparing Perplexity against the other two classes of methods, we select the best-performing method as representative [^2] For succinctness, we adopt this measure throughout the paper.
Perplexity does not perform as well as GPTSniffer, which is a pre-training-based method in all eight programming languages. The AUC of Perplexity and GPTSniffer for each language are as follows: C (90.40%–99.13%), C# (90.12%–98.15%), C++ (91.37%–99.58%), Go (72.68%–94.19%), Java (85.54%–96.11%), Python (63.89%–95.17%), Ruby (73.44%–93.25%), and Rust (89.72%–98.64%). The difference can be explained that GPTSniffer is a supervised method while Perplexity is not. Moreover, Perplexity cannot achieve the same performance as feature-based methods in Python and C: Python (63.89%–89.60%), C (90.40%–99.01%).
When comparing the AUC of Perplexity between programming languages, we observe that Perplexity performs best in C/C++ in which NPR and DetectCodeGPT achieve an average AUC of 85%. However, NPR, DetectCodeGPT, and other methods cannot achieve an AUC of over 80% in other languages. Taking C/C++ and Python as examples. C/C++ requires developers to implement many functions, which are more likely to be provided by third-party libraries in Python. In this case, C/C++ has more code stylometry than Python (Caliskan-Islam et al., 2015). However, the design philosophy of Python encourages simplified code stylometry, making it easy for LLMs to generate human-style code.
$\textbf{RQ}_{\textbf{1.1}}$: Perplexity can not perform as well as the feature- and pre-training-based methods from the perspective of programming languages. Among eight languages, Perplexity is only effective in C and C++.
5.2. $\text{RQ}_{\text{1.2}}$: Degree of Difficulty
Table 4 and Table 5 present the AUC of the three groups of methods from the perspective of the degree of difficulty.
Table 4. The AUC of Perplexity in terms of degrees of difficulty
Difficulty Code LLM without perturbation with perturbation
@@
Table 5. The AUC of feature- and pre-training-based methods in terms of degrees of difficulty
Difficulty Feature Pre-training
@@
Perplexity exhibits lower performance compared with the feature- and pre-training-based methods in terms of all three degrees of difficulty. The differences between Perplexity and pre-training-based methods for each level of difficulty are as follows: easy (75.41%–90.57%), middle (77.00%–94.32%), and hard (76.48%–94.33%). When compared with feature-based methods, Perplexity is less effective than Idialu: easy (75.41%–85.37%), middle (77.00%–89.60%), and hard (76.48%–93.18%), and Bukhari: easy (75.41%–96.40%), middle (77.00%–98.98%), and hard (76.48%–98.25%). The discrepancy can be explained that feature- and pre-training-based methods are based on supervised training, while Perplexity not.
When it comes to the AUC of Perplexity from degrees of difficulty, we find that Perplexity does not perform well in all three degrees of difficulty, but slightly improves from the easy to the hard level. To be specific, log p(x) and Log Rank are the best-performing methods in Perplexity without perturbation, achieving the AUC of 61.25% and 60.10% at the easy level, 62.00%, and 61.76% at the middle level, and 63.54% and 64.53% at the hard level, respectively. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve the AUC of 66.40% and 65.60% at the easy level, 67.82% and 67.67% at the middle level, and 68.95% and 69.64% at the hard level, respectively. Since the solutions to difficult puzzles have complex logic and customized implementations, they have more code stylometry than solutions to easy puzzles.
$\textbf{RQ}_{\textbf{1.2}}$: Perplexity perform not well in all three degrees of difficulty, even though the performance slightly improves from the easy to the hard level. The feature- and pre-training-based methods perform well at all three degrees of difficulty.
5.3. $\text{RQ}_{\text{1.3}}$: Scale of Solution
Table 6 and Table 7 show the AUC of three classes of methods in terms of the scale of solutions. Note that in Table 7, we provide results only in the small and middle scales for Idialu due to no large-scale code solution in Python.
Table 6. The AUC of Perplexity in terms of scales of solution
@@
Table 7. The AUC of feature- and pre-training-based methods in terms of scales of solution
@@
Perplexity does not perform as well as the pre-training-based method, i.e., GPTSniffer in all three scales. The AUC of Perplexity and GPTSniffer for each scale is: small (69.75%–82.74%), middle (77.42%–96.10%), and large (87.81%–94.19%). For the featured-based method, Perplexity is less effective than Idialu: small (69.75%–85.87%), middle (77.42%–92.58%), and Bukhari: small (69.75%–91.20%), middle (77.42%–98.26%), and large (87.81%–96.67%). These differences are mainly because feature- and pre-training-based methods are supervised, while Perplexity is a collection of heuristic methods.
@@
When comparing the methods in Perplexity regarding different scales, we observe that those well-performed methods improve from the small to the large scale. For example, log p(x) is the best-performing method in Perplexity without perturbation, with an AUC of 61.41% for small, 66.97% for middle, and 77.53% for large. Similarly, the best-performing methods in Perplexity with perturbation, i.e., NPR and DetectCodeGPT, achieve an AUC of 64.87% and 63.06% for small, 72.91% and 72.19% for middle, and 76.81% and 78.10% for large, respectively. These results can be explained that large-scale code snippets have more diverse and detailed implementations than others, which makes HaCode more distinct from LLMgCode.
@@
$\textbf{RQ}_{\textbf{1.3}}$: Perplexity performs less effectively compared with feature- and pre-training-based methods in all the three scales of code solutions. Perplexity only exhibits acceptable performance in the large-scale code solution.
5.4. $\text{RQ}_{\text{2}}$: Detection Speed